### v1

# Qffusion: Controllable Portrait Video Editing via Quadrant-Grid Attention Learning

Abstract
This paper presents Qffusion, a dual-frame-guided framework for portrait video editing. Specifically, we consider a design principle of “animation for editing”, and train Qffusion as a general animation framework from two still reference images while we can use it for portrait video editing easily by applying modified start and end frames as references during inference. Leveraging the powerful generative power of Stable Diffusion, we propose a Quadrant-grid Arrangement (QGA) scheme for latent re-arrangement, which arranges the latent codes of two reference images and that of four facial conditions into a four-grid fashion, separately. Then, we fuse features of these two modalities and use self-attention for both appearance and temporal learning, where representations at different times are jointly modeled under QGA. Our Qffusion can achieve stable video editing without additional networks or complex training stages, where only the input format of Stable Diffusion is modified. Further, we propose a Quadrant-grid Propagation (QGP) inference strategy, which enjoys a unique advantage on stable arbitrary-length video generation by processing reference and condition frames recursively. Through extensive experiments, Qffusion consistently outperforms state-of-the-art techniques on portrait video editing.
Figure 1: We present Qffusion, a simple yet effective dual-frame-guided portrait video editing framework. Specifically, our Qffusion is trained as a general animation framework from two still reference images whereas it can perform portrait video editing effortlessly when using modified start and end video frames as references during inference. That is, we specify editing requirements by modifying two video frames rather than text. Our Qffusion can perform fine-grained local editing (e.g., modifying age, makeup, hair, style, and sunglasses).
footnotetext: * Eual contribution, $\dagger$ Corresponding authors.
1 Introduction
With the rapid proliferation of mobile internet and video platforms, portrait video editing has become one of the cornerstones of computer graphics and vision. Traditional approaches often require professional designers with time-consuming processes, e.g., scene setup, staged recording, editing, and repetitive iterations, which are laborious and inefficient. Considering fruitful endeavors have been pursued in image generation [43, 45, 44, 60] and video generation [5, 1, 8, 2] in recent years, employing powerful large-scale generative models to assist portrait video editing has become feasible.
Generally, the techniques for portrait video editing can be classified into two categories. One category is based on Generative Adversarial Networks (GANs) [17, 40]. While these methods provide decent processing speeds, they often fail to generalize to unseen humans and suffer from unstable training [54]. The other category comprises diffusion-based methods [20, 47]. These models [53, 10, 41, 16, 28, 27] are pretrained on the large image and video datasets, making them easily model arbitrary objects. Despite the generalizability and controllability of these diffusion models, they still face several challenges, which are listed in Tab. 1. (i) The prevailing text-driven video editing methods[53, 10, 41, 16, 28] struggle to deal with specific local manipulations (e.g., hair editing) since text-driven editing cannot capture sufficient editing details directly. (ii) The very recent method AnyV2V [27] performs first-frame-guided editing, which first employs an image editing model [7, 59] to modify the first frame and then utilizes an image-to-video (I2V) generation model  [62] to propagate such modifications. However, it often leads to a degraded quality, as a single edited frame cannot enforce sufficient appearance coherence. (iii) The video length of the existing methods is always constrained by limited computational resources.
To handle the above challenges, this paper proposes a dual-frame-guided portrait video editing method dubbed Qffusion, which allows for fine-grained local editing on arbitrary-long videos. Specifically, we consider an “animation for editing” design principle and train Qffusion as a video animation network from two still reference images, which can perform portrait video editing effortlessly when applying edited start and end video frames as references during inference. That is, our Qffusion specifies editing requirements by modifying two video frames rather than text or one single frame. For example, we can use professional software (e.g., Photoshop or Meitu11https://www.meitu.com/) for controllable and consistent reference frame editing.
Settings Fine-grained Arbitrary-long Non-degraded appearance
Text
First-frame
Dual-frame
Table 1: Categories and characteristics of diffusion-based video editing methods. This paper is the first to propose dual-frame settings for high-quality portrait video editing.
Concretely, we design a Quadrant-grid Arrangement (QGA) scheme into image models (i.e., Stable Diffusion [44]) for video modeling, which only modifies the input format for training. Here, a four-grid representation is designed for two reference images and four sequential driving keypoints, respectively. In detail, we organize two reference images and two all-zero placeholders as intermediate masks into a big four-grid image. Then, this four-grid image representation is stacked with the corresponding four-grid driving representation. Benefiting from the feature aggregation ability of the attention mechanism, QGA scheme can establish the correspondence between driving conditions and reference appearance, where temporal clues are also modeled as motion information is embraced naturally in the four-grid driving representation. Moreover, to make an even motion modeling during inference, we design a Quadrant-grid Propagation (QGP) inference algorithm, which recursively uses generated frames at the current inference iteration as reference frames for the next iteration, making the edited video length unconstrained.
As shown in Fig. 1, Qffusion delivers impressive results in fine-grained local editing, e.g., adding sunglasses, editing age, hair, and style. Besides, since Qffusion is trained as a general video animation framework, we can flexiblely use it for other applications, e.g., whole-body driving [22] and jump cut smooth [51]. Our main contributions are:
  • We propose a novel dual-frame-guided framework for portrait video editing, which propagates fine-grained local modification from the start and end video frames.
  • We propose a Quadrant-grid Arrangement (QGA) scheme to re-arrange reference images and driving signals under a four-grid fashion respectively, which models appearance correspondence and temporal clues all at once.
  • We propose a recursive inference strategy named Quadrant-grid Propagation (QGP), which can stably generate arbitrary-long videos.
  • Our Qffusion can deliver rich application extensions, e.g., portrait video editing, whole-body driving [22], and jump cut smooth [51], showing more competitive results with those state-of-the-art task-specific methods.
2 Related Works
Diffusion Model for Image Generation and Editing. Recently, diffusion models [20, 47] have emerged as a popular paradigm for text-to-image (T2I). DALLE-2 [43] and Imagen [45] can generate high-resolution images via cascaded diffusion models. Then, Stable Diffusion [44] proposes to train diffusion models in the learned latent space for less computational complexity. As for image editing, the early techniques [15, 38, 43, 3, 36] need an editing mask provided by the user, which is time-consuming. To deal with this, there is a line of research conducts text-only image editing [19, 49, 13, 37, 9, 4, 35], which changes the visual content of the input image following the target prompt without masks. Moreover, InstructPix2Pix [7] and MagicBrush [59] perform editing following human instructions. To further improve spatial controllability, ControlNet [60] introduces a side path to Stable Diffusion to accept extra conditions like edges, depth, and human pose.
Diffusion Model for Video Generation and Editing. There have also been substantial efforts in Text-to-Video (T2V). Besides operating diffusion process directly on pixel space [21, 46], the recent T2V models [6, 64, 14, 50] draw inspiration from Stable Diffusion [44] and generate high-quality videos via a learned latent space. Apart from T2V, several representative works [5, 1, 8, 2, 62] lay the cornerstone for image-to-video (I2V) generation.
Regarding text-driven video editing, Tune-A-Video [53] first proposes an efficient one-shot tuning strategy based on Stable Diffusion. Then, a group of methods [16, 28, 10, 41, 11, 56] conduct zero-shot video editing, where various attention mechanisms are designed to capture temporal cues without extra training. Further, following the substantial efforts of I2V generation, AnyV2V [27] conducts image-driven video editing by propagating modified content from the first edited frame. However, it leads to a degraded quality, where those frames that are far away from the first frame usually present an unpleasing reconstruction and editing.
Generally, text-driven video editing methods struggle with certain fine-grained local manipulations on portrait videos, e.g., modifying hair. Different from the typical text-driven video editing methods, Codef [39] introduces a new type of video representation, which consists of a canonical content field and a temporal deformation field recording static contents and transformations separately. By editing the canonical image, Codef can carry out fine-grained local editing. However, it needs training on each video to be edited, whereas ours is a general framework after training.
Diffusion-based Video Animation. In recent years, apart from pose-controllable text-to-video generation [34], some researchers focus more on generating animated videos from still images with diffusion models. DreamPose [25] proposes a two-stage finetuning strategy with pose sequence. BDMM [57] designs a Deformable Motion Modulation that utilizes geometric kernel offset with adaptive weight modulation for subtle appearance transfer. Besides, Animate Anyone [22] temporally maintains consistency by a ReferenceNet merging detail features via spatial attention, where a pose-guided module is designed for movements. Unlike them, our Qffusion is very flexible for various applications, such as portrait video editing, whole-body driving [22], and jump cut smooth [51].
3 Preliminary of Stable Diffusion
Stable Diffusion [44] consists of a VAE [26], a diffusion process, and a denoising process. Here, VAE provides a learnable latent space, avoiding the massive resources required for pixel-level calculation.
Diffusion Process. In the diffusion process, the model progressively corrupts input data $\mathbf{z}_{0}\sim p(\mathbf{z}_{0})$ according to a predefined schedule $\beta_{t}\in(0,1)$, turning data distribution into an isotropic Gaussian in $T$ steps. Formally, it can be expressed as:
$q\left(\mathbf{z}_{1:T}\mid\mathbf{z}_{0}\right)$ $=\prod q\left(\mathbf{z}_{t}\mid\mathbf{z}_{t-1}\right),\qquad{t\in[1,...,T]}.$ (1)
Denoising Process. In the denoising process, the model learns to invert the diffusion procedure so that it can turn noise into real data distribution at inference. The corresponding backward process can be described as follows:
$p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t})$ $=\mathcal{N}(\mu_{\theta}(\mathbf{z}_{t},t),\Sigma_{\theta}(\mathbf{z}_{t},t))$ (2)
$=\mathcal{N}(\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{z}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\epsilon}),\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}),$
[…]
$\mathcal{L}_{t}$ $\!=\!\mathbb{E}_{\mathbf{z}_{0},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\left\|{\epsilon}\!-\!\mathbf{\epsilon}_{\theta}\left(\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}\!+\!\sqrt{1-\bar{\alpha}_{t}}{\epsilon};t\right)\right\|^{2}\right].$ (3)
To make the model conditioned on extra condition $\mathbf{z}_{c}$, we can inject $\mathbf{c}$ into $\mathbf{\epsilon}_{\theta}(\cdot)$ by replacing $\mu_{\theta}\left(\mathbf{z}_{t},t\right)$ and $\Sigma_{\theta}\left(\mathbf{z}_{t},t\right)$ with $\mu_{\theta}\left(\mathbf{z}_{t},t,\mathbf{c}\right)$ and $\Sigma_{\theta}\left(\mathbf{z}_{t},t,\mathbf{c}\right)$.
Figure 2: Overview illustration of Qffusion. As for training, we first design a Quadrant-grid Arrangement (QGA) scheme for latent re-arrangement, which arranges the latent codes of two reference images and that of four portrait landmarks into a four-grid fashion, separately. Then, we fuse features of these two modalities and use self-attention for both appearance and temporal learning. Here, the facial identity features [12] are also put into cross-attention mechanism in the denoising U-Net for further identity constraint. During inference, a stable video is generated via our proposed Quadrant-grid Propagation (QGP) strategy.
4 Methods
We propose a dual-frame-guided portrait video editing method dubbed Qffusion, which can perform fine-grained or local editing on arbitrary-long videos. Specifically, we consider an “animation for editing” principle, and train Qffusion as a video animation framework from two still reference images while we can use it for portrait video editing easily by applying edited start and end video frames as references during inference. That is, we first specify editing requirements by modifying the start and end frames with professional software and then use Qffusion to propagate these fine-grained local modifications to the entire video.
The Section is organized as follows: Sec. 4.1 first introduces an overview of our proposed Qffusion. Sec. 4.2 illustrates the Quadrant-grid Arrangement (QGA) scheme in SD for latent re-arrangement. Then, our recursive inference strategy Quadrant-grid Propagation (QGP) for stable and arbitrary-length video generation is presented in Sec. 4.3.
4.1 Overview
The overall pipeline of our Qffusion is illustrated in Fig. 2. Like SD, our model consists of two parts, i.e., VAE and latent diffusion. In Qffusion, we propose a QGA scheme to arrange four sequential frames into a large four-grid image, where the upper-right and bottom-left frames are masked for generation. This four-grid image is then stacked with their corresponding four-grid keypoints, forming a composite input that encodes both visual and motion information. Formally, given four sequential frames $\{\mathbf{I}^{a},\mathbf{I}^{b},\mathbf{I}^{c},\mathbf{I}^{d}\}$ and their corresponding condition image (i.e., keypoints) $\{\mathbf{C}^{a},\mathbf{C}^{b},\mathbf{C}^{c},\mathbf{C}^{d}\}$, we replace $\mathbf{I}^{b}$ and $\mathbf{I}^{c}$ with all-zero masks and train our model to reconstruct them. The VAE encoder $\mathcal{E}$ first encodes $\mathbf{I}^{a},\mathbf{I}^{d},\mathbf{C}^{a},\mathbf{C}^{b},\mathbf{C}^{c},\mathbf{C}^{d}$ into latent codes $\mathbf{r}^{a},\mathbf{r}^{d},\mathbf{c}^{a},\mathbf{c}^{b},\mathbf{c}^{c},\mathbf{c}^{c}$, respectively. Then, these input latent codes are combined with a noise map to form a fused code through our QGA scheme. Next, a denoiser $\mathbf{\epsilon}_{\theta}$ learns driving correspondence and temporal clues from these latent codes and predicts the denoised latent. Finally, a VAE decoder $\mathcal{D}$ decodes the denoised latent into images. The computation process of Qffusion is:
$\tilde{\mathbf{I}}^{b},\tilde{\mathbf{I}}^{c}=\text{Qffusion}(\mathbf{I}^{a},\mathbf{I}^{d},\mathbf{C}^{a},\mathbf{C}^{b},\mathbf{C}^{c},\mathbf{C}^{d}).$ (4)
In summary, Qffusion takes two frames $\mathbf{I}^{a}$ and an $\mathbf{I}^{d}$ as appearance references, where the condition images of references $\mathbf{C}^{a}$ and $\mathbf{C}^{d}$ and that of intermediate frames ($\mathbf{C}^{b}$ and $\mathbf{C}^{c}$) as motion signals for the generation of $\tilde{\mathbf{I}}^{b}$ and $\tilde{\mathbf{I}}^{c}$. After training, the model would generate two portrait frames at each inference time. By replacing the intermediate conditions sequentially, our method can generate arbitrary-length videos easily.
4.2 Quadrant-grid Arrangement
Based on SD, we train Qffusion as a general video animation framework from two reference images and four driving signals. Specifically, we propose a Quadrant-grid Arrangement (QGA) scheme to establish the correspondence between two modalities (i.e., appearance features and driving signals) for appearance consistency in the denoiser UNet $\mathbf{\epsilon}_{\theta}$, while jointly modeling the temporal clues at different times.
QGA arranges the latent codes of reference images ($\mathbf{r}^{a}$ and $\mathbf{r}^{d}$) and two all-zero placeholder masks for intermediate frames into a big four-grid image $\{\mathbf{r}^{a},\mathbf{0},\mathbf{0},\mathbf{r}^{d}\}$, where reference images are assigned to upper-left and bottom-right locations. Similarly, we combine four driving conditions into a big four-grid condition image as $\{\mathbf{c}^{a}$, $\mathbf{c}^{b}$, $\mathbf{c}^{c},\mathbf{c}^{d}\}$, which is stacked with the previous four-grid appearance latents $\{\mathbf{r}^{a},\mathbf{0},\mathbf{0},\mathbf{r}^{d}\}$. Here, a one-to-one correspondence between the appearance and conditions is achieved. In this way, appearance representations of different frames would establish spatial relationships in the self-attention layers for the reconstruction of $\tilde{\mathbf{I}}^{b}$ and $\tilde{\mathbf{I}}^{c}$. Besides, temporal clues are also modeled naturally in QGA since motion information is embraced in the composed four-grid representation.
As illustrated in Fig. 2, QGA scheme stacks the four-grid representations of reference frames ($\mathcal{R}$) and that of condition images ($\mathcal{C}$), and then combines a same-size noise map ($\mathcal{Z}_{t}$), thus obtaining the fused latent representation $\mathcal{Q}_{t}$ as:
$\mathcal{Q}_{t}$ $=\texttt{QGA}(\mathbf{I}^{a},\mathbf{C}^{a},\mathbf{C}^{b},\mathbf{C}^{c},\mathbf{C}^{d},\mathbf{I}^{d},t)=\mathcal{R}\odot\mathcal{C}\odot\mathcal{Z}_{t}$ (5)
$=\left[\begin{array}[]{cc}\mathbf{r}^{a}&\mathbf{0}\\ \mathbf{0}&\mathbf{r}^{d}\end{array}\right]\odot\left[\begin{array}[]{cc}\mathbf{c}^{a}&\mathbf{c}^{b}\\ \mathbf{c}^{c}&\mathbf{c}^{d}\end{array}\right]\odot\left[\begin{array}[]{cc}\mathbf{z}_{t}^{a}&\mathbf{z}_{t}^{b}\\ \mathbf{z}_{t}^{c}&\mathbf{z}_{t}^{d}\end{array}\right],$
where $\mathbf{c}^{b}$ and $\mathbf{c}^{c}$ are driving latent codes from intermediate condition images, i.e., $\mathbf{c}^{b}=\mathcal{E}(\mathbf{C}^{b})$ and $\mathbf{c}^{c}=\mathcal{E}(\mathbf{C}^{c})$. $\mathbf{z}_{t}^{*}$ is the noise of *-th frame at timestep $t$. $\odot$ denotes channel-wise concatenation. Each latent code in the inputs of QGA (i.e., $\mathbf{r}^{*},\mathbf{c}^{*},\mathbf{z}^{*}$) is with the size of $\mathbb{R}^{B\times C\times H\times W}$ and thus $\mathcal{Q}_{t}\in\mathbb{R}^{B\times 3C\times 2H\times 2W}$, where $B$ denotes batch size.
Next, the fused $\mathcal{Q}_{t}$ is fed to denoiser $\mathbf{\epsilon}_{\theta}$ to predict the denoised latent codes. During training, the diffusion process iteratively adds noises to $\mathcal{Z}_{0}$ and eventually leads to $\mathcal{Z}_{T}$. In the denoising process, the denoiser $\mathbf{\epsilon}_{\theta}$ aims to recover latent codes $[[\mathbf{z}^{a},\mathbf{z}^{b}]^{\top},[\mathbf{z}^{c},\mathbf{z}^{d}]^{\top}]$ based on the fused latent $\mathcal{Q}_{T}$.
Utilizing the learning framework in SD, our method only adjusts the number of I/O channels of UNet. After the denoising process, we can obtain the generated intermediate frame $\tilde{\mathbf{I}}^{i}=\mathcal{D}(\tilde{\mathbf{z}}^{i})$, where $i\in\{b,c\}$ and $\tilde{\mathbf{z}}^{i}$ is generated by splitting and unstacking the denoised $\tilde{\mathcal{Q}}_{0}=\epsilon_{\theta}(\tilde{\mathcal{Q}}_{1}),\dots,\tilde{\mathcal{Q}}_{T-1}=\epsilon_{\theta}(\mathcal{Q}_{T})$:
$\tilde{\mathbf{z}}^{i}=\tilde{\mathcal{Z}}_{0}[:,:,:H,W:],~{}~{}s.t.~{}~{}\tilde{\mathcal{Z}}_{0}=\tilde{\mathcal{Q}}_{0}[:,2C:,:].$ (6)
In this way, our training objective can be expressed as:
$\mathcal{L}_{t}^{\prime}\!=\!\mathbb{E}_{\mathcal{Z}_{0},\mathcal{R},\mathcal{C},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\!\!\left[\!\left\|{\epsilon}\!-\!\mathbf{\epsilon}_{\theta}\!\!\left(\sqrt{\bar{\alpha}_{t}}\mathcal{Z}_{0}\!+\!\sqrt{1-\bar{\alpha}_{t}}{\epsilon};t,\mathcal{R},\mathcal{C}\right)\!\right\|^{2}\!\right].$
4.3 Quadrant-grid Propagation for Inference
Figure 3: The illustration of Quadrant-grid Propagation (QGP) for stable arbitrary-length video generation. At each iteration, we use the same intervals between each sub-images in the four-grid representation to bring even temporal modeling. We omit the rounding symbols for clarity.
The remaining problem is how to continuously generate all intermediate frames given the start and end reference frames and driving signals during inference. Considering our quadrant-grid training design, we also maintain this fashion for inference. Specifically, we only generate two frames for each inference time. Then, we can generate arbitrary-length videos easily by replacing the intermediate conditions sequentially. This makes portrait video editing no longer constrained by limited computing resources.
Naive Inference. We assume an input portrait video with $K+1$ frames, which are indexed are $[0,1,2,...K]$ respectively. A naive inference way is fixing two reference images as 0-th and $K$-th frames to gradually generate intermediate frames, ie, $\{1,K-1\},\{2,K-2\},...,\{\left\lfloor{K/2}\right\rfloor,\left\lfloor{(K+1)/2}\right\rfloor\}$. This requires $K/2$ inference times. However, the synthesized frames suffer from the issue of excessive interval: (1) the interval between two generated frames is excessive, e.g., a $K-2$ gap between the synthesized 1-th and $(K-1)$-th frame, (2) the intervals between intermediate frames and reference frames are excessive, e.g., a $\left\lfloor{K/2}\right\rfloor$ gap between the synthesized $\left\lfloor{K/2}\right\rfloor$ frame and 00-th reference frame. The issue would lead to unstable motion modeling, especially when $K$ is big, making the naive inference a suboptimal solution.
QGP Inference. In order to seek a more even motion modeling, this paper proposes a recursive influence strategy Quadrant-grid Propagation (QGP), in which we use generated frames at the current inference iteration as reference frames for the next iteration. As shown in Fig. 3, we first use 0-th and $K$-th reference frames to generate the intermediate two frames, which are indexed as $\left\lfloor\frac{1}{3}K\right\rfloor$-th and $\left\lfloor\frac{2}{3}K\right\rfloor$-th. Then in the next iteration, the newly generated $\left\lfloor\frac{1}{3}K\right\rfloor$-th and $\left\lfloor\frac{2}{3}K\right\rfloor$-th frames would serve as reference frames for their corresponding intermediate frames. The process is stopped until all $K+1$ frames are synthesized. Specifically, the quadrant arrangement at each inference iteration would be: $\left\{\left[0,\left\lfloor\frac{K}{3}\right\rfloor,\left\lfloor\frac{2K}{3}\right\rfloor,K\right]\right\},\{\left[0,\left\lfloor\frac{K}{9}\right\rfloor,\left\lfloor\frac{2K}{9}\right\rfloor,\left\lfloor\frac{K}{3}\right\rfloor\right]$, $\left[\left\lfloor\frac{K}{3}\right\rfloor,\left\lfloor\frac{4K}{9}\right\rfloor,\left\lfloor\frac{5K}{9}\right\rfloor,\left\lfloor\frac{2K}{3}\right\rfloor\right],\left[\left\lfloor\frac{2K}{3}\right\rfloor,\left\lfloor\frac{7K}{9}\right\rfloor,\left\lfloor\frac{8K}{9}\right\rfloor,K\right]\,\}$, ,$,\{[0,1,2,3],...\}$. Note that at each iteration, QGP makes the intervals between each sub-image in the four-grid representation the same, leading to a more even temporal sampling than naive inference.
Summing up, although both the proposed QGP and the naive inference can synthesize arbitrary-length videos, the former can carry out smoother temporal modeling by making the interval of generated frames the same. More implementation details can be found in our Appendix.
Figure 4: Animation comparisons with other methods.
5 Experiment
5.1 Implementation Details
Training. We use a fixed interval of 5 on video sequences to collect four frames as $\{\mathbf{I}^{a},\mathbf{I}^{b},\mathbf{I}^{c},\mathbf{I}^{d}\}$ in the QGA scheme. Based on SD 1.5, Qffusion uses an AdamW optimizer [31] with gradient accumulations set to 2. For the learning rate schedule, a warm-up strategy is applied, which gradually increases the learning rate to 0.0001 throughout 10,000 steps. Other hyper-parameters are followed with SD. We train all of our models on an NVIDIA A100 GPU, where a VAE and a denoiser $\epsilon_{\theta}$ are trained, respectively. For VAE training, the batch size is set to 4. The VAE training takes about 8 hours. For denoiser training, the batch size is set to 1. The denoiser training takes about 8 hours.
Experimental setup. We train Qffusion on HDTF dataset [63] and evaluate it on LSP [32] and some videos in RAVDESS [30] and Celebv-HQ [65] datasets. Each video contains a high-resolution portrait. The average video length is 1-5 minutes processed at 25 fps. Each video is cropped to keep the face at the center and resized to $256\times 256$. LSP contains 4 video sequences. Our conditions consist of dense facial landmarks and torso lines. We detect 478 3D facial landmarks for all videos using Mediapipe [33]. The 3D torso points describing the shoulder boundaries are estimated by [29]. Besides, we use professional software (e.g., Photoshop or Meitu) to edit start and end frames to maintain consistency.
Evaluation Metrics. To verify the effectiveness of our Qffusion on portrait video animation, we use the average Peak signal-to-noise Noise Ratio (PSNR) [24], Structural Similarity Index Measure (SSIM) [52], and Learned perceptual similarity (LPIPS) [61]. Besides, we apply Warp Error [10, 16] to measure the temporal consistency of generated videos. Specifically, we first estimate optical flow [48] of the input video and then use it to warp the generated frames. Next, the average MSE between each warped frame and the target ones is calculated.
[…]
(a) GAN 17.64 0.548 0.654 0.350
(b) ControlNet 18.29 0.544 2.124 0.322
(c) ControlNet+AnimateDiff 10.69 0.212 20.28 0.636
(d) Ours 25.22 0.834 0.665 0.154
Table 2: Quantitative comparisons with state-of-the-art animation methods, where our method yields the best performance on PSNR, SSIM, LPIPS, and competitive Warp Error.
To evaluate portrait video editing, we first use CLIP-Image similarity to measure the reference alignment of edited videos. It computes the average cosine similarity of image embeddings from CLIP model [42] between the edited video frame and the rest of the generated frames. Warp Error is also leveraged to measure the temporal consistency of edited videos.
5.2 Qualitative Comparison on Animation
Since Qffusion is trained as a video animation framework, we first present a visual comparison with three portrait video animation methods in Fig. 4, which are: (i) GAN. A UNet-based GAN [32], which is trained for reconstruction. Here, two reference images and condition images are used as input to predict corresponding animation images. (ii) ControlNet. We apply ControlNet [60] to encode the condition images and the reference ones. Specifically, we first use two ”reference-only” ControlNet2 2https://github.com/Mikubill/sd-webui-controlnet/discussions/1236 to encode two reference images, and an ”OpenPose” ControlNet3  3https://huggingface.co/lllyasviel/sd-controlnet-openpose for condition encoding. Then, all encoded features are fed into SD to generate animated images that follow the condition motion and reference appearance. (iii) ControlNet $+$ AnimateDiff. We insert the temporal module of AnimateDiff [18] into ControlNet for temporal consistency.
Although Qffusion can generate arbitrary-length videos, we use all methods to generate 80 frames for comparison here. Fig. 4 shows the animation performance, where two reference images are omitted for simplicity. Our Qffusion has the most consistent appearance details and motion. Note that AnimateDiff only generates 16-32 frames. To generate 80 frames, we use an overlap generation strategy (i.e., overlapping 8 frames for a 16-frame generation) to maintain continuity.
Figure 5: Qualitative ablation studies of input arrangement. (a) two images $\{\mathbf{r}^{a},\mathbf{0}\}$; (b) a four-grid arrangement with one reference image $\{\mathbf{r}^{a},\mathbf{0},\mathbf{0},\mathbf{0}\}$ ; (c) our quadrant-grid design $\{\mathbf{r}^{a},\mathbf{0},\mathbf{0},\mathbf{r}^{d}\}$; (d) ground-truth, respectively.
[…]
(a) $\{\mathbf{r^{a}},\mathbf{0}\}$ 19.78 0.648 1.985 0.272
(b) $\{\mathbf{r^{a}},\mathbf{0},\mathbf{0},\mathbf{0}\}$ 23.87 0.803 1.760 0.165
(c) Our $\{\mathbf{r^{a}},\mathbf{0},\mathbf{0},\mathbf{r^{d}}\}$ 25.22 0.834 0.665 0.154
Table 3: Quantitative ablation studies on different settings of input arrangement.
5.3 Quantitative Comparison on Animation
We report PSNR, SSIM, Warp Error, and LPIPS for quantitative comparison in Table 2. Our Qffusion excels the current state-of-the-art methods by a large margin on PSNR, SSIM, and LPIPS. Although GAN can achieve a slightly superior Warp Error than our Qffusion, it yields the worst ID fidelity in Fig. 4. Besides, Qffusion achieves the best Warp Error among diffusion-based methods, which shows our capacity for long-term temporal modeling by Quadrant-grid Arrangement (QGA) and Quadrant-grid Propagation (QGP). Besides, although the temporal module of AnimateDiff can bring temporal prior, it only supports fixed and limited length generation. Even if the overlap strategy is used for long video generation, the Warp Error of ControlNet $+$ AnimateDiff is still the worst.
Figure 6: Qualitative ablation of our QGP inference and the naive inference, where QGP results are more motion-aligned and appearance-consistent.
[…]
Naive Inference 25.01 0.827 2.488 0.160
Our QGP 25.22 0.834 0.665 0.154
Table 4: Quantitative ablation studies of naive inference strategy and the proposed recursive QGP inference strategy.
5.4 Ablation Studies
Figure 7: We compare with three recent video editing methods: TokenFlow [16], Codef [39] and AnyV2V [27] on portrait video editing. TokenFlow cannot deal with certain local editing (e.g., hair editing), Codef always has artifacts, and AnyV2V struggles for a pleasing appearance. Our method requires both modified start and end frames ($\mathbf{I}^{s}$ and $\mathbf{I}^{e}$) as editing signals, where $\mathbf{I}^{e}$ is omitted here for simplicity.
Figure 8: Our Qffusion can realize fine-grained local editing, including changing age and style, and adding beauty masks. Note that Qffusion needs the modified start and end frame ($\mathbf{I}^{s}$ and $\mathbf{I}^{e}$) for editing propagation, where $\mathbf{I}^{e}$ is omitted for simplicity.
Figure 9: Whole-body driving results. We compare our Qffusion with DreamPose [25], BDMM [57], Animate Anyone [22]. We achieve better results than those task-specific methods, where the end reference is omitted for simplicity.
Quadrant-grid Arrangement. We conduct ablation studies to validate the effectiveness of our key design of QGA. For a fair comparison, we use the same training and inference strategies. We report the performance of the following three settings for latent arrangement. As illustrated in Tab. 3, (a) Two images side-by-side ($\{\mathbf{r}^{a},\mathbf{0}\}$): a reference image is concatenated with an all-zero placeholder mask side by side. (b) Four-grid with one reference image ($\{\mathbf{r}^{a},\mathbf{0},\mathbf{0},\mathbf{0}\}$): a reference image is arranged in the left-top corner of a four-square grid, leaving the remaining three squares to be zeros. (c) The proposed quadrant-grid design QGA ($\{\mathbf{r}^{a},\mathbf{0},\mathbf{0},\mathbf{r}^{d}\}$).
Quantitative results of our quadrant-grid design are shown in Tab. 3. Both (a) and (b) use only one reference image as a start point for generation. However, (a) and (b) yield inferior performance in temporal consistency and image quality for portrait video animation. Our QGA, on the contrary, uses two reference frames to constrain the generation of intermediate frames, which achieves significant gain over (a) and (b), demonstrating our effectiveness.
We also perform a qualitative evaluation to verify the effectiveness of our quadrant-grid design in Fig. 5. The results of (a) exhibit noticeable lighting jitter ($1$-st row, $2$-nd column) and severe artifacts (in $1$-st row, $3$-rd and $4$-th column). The results of (b) show color and lighting jitters among frames ($2$-nd row, $2$-nd column), and inaccurate mouth movement ($2$-nd row, $4$-th column). As shown in the $3$-rd row, our QGA can generate temporal-consistent portrait videos. To sum up, compared with one reference, we argue that two references can help regularize temporal appearance in the generated video.
Quadrant-grid Propagation. To validate the effectiveness of our QGP inference, we compare it with the naive inference in Tab. 4. QGP outperforms the naive inference on all metrics, especially on Warp Error, which decreases 72.6% dramatically. The excellent Warp Error of QGP shows its superiority in long-term temporal consistency, which achieves intervals between each sub-image more even in our quadrant-grid design. Besides, we provide the qualitative comparison between the proposed QGP inference and the naive inference in Fig. 6, which also support our findings in Tab. 4.
5.5 Applications
Our Qffusion can deliver three applications: portrait video editing, whole-body driving [22], and jump cut smooth [51].
Portrait video editing. We compare Qffusion with the current state-of-the-art video editing methods Codef [39], TokenFlow [16] and AnyV2V [27] in Fig. 7. The editing scenarios consist of modifying style and hair, and adding sunglasses. As a text-driven editing method, TokenFlow [16] cannot deal with some fine-grained local editing, such as hair editing. Codef [39] sometimes suffers from the inconsistent ID issue (left example). Note that Codef [39] relies on generating a canonical image to record static content in a video, which however usually are with artifacts. When we further edit it with the desired style or appearance, these artifacts will also be spread throughout the entire video. In contrast, our editing results are identity-consistent and abide by the conditional poses clearly. Moreover, although AnyV2V [27] performs image-driven video editing, it faces the degradation of editing appearance. Further, none of the existing methods can generate arbitrary-long videos.
We also provide more examples of our Qffusion on portrait video editing in Fig. 8, where more intermediate frames are given. To further evaluate the video editing ability of Qffusion, we provide a quantitative comparison in Tab. 5. Specifically, we report the average CLIP-Image similarity and Warp Error. Our method yields the best performance on CLIP score and Warp Error, which demonstrates that Qffusion can achieve amazing reference alignment and motion consistency. Note that we cannot calculate CLIP-Image for TokenFlow [16], since it is a text-driven method. Besides, one may wonder whether we can use image editing tools [7, 59] for portrait video editing. We argue that per-frame editing often fails to preserve temporal consistency. Thus, we do not compare with per-frame editing results.
[…]
TokenFlow [16] - 0.874
Codef [39] 0.938 0.301
AnyV2V [27] 0.924 0.822
Ours 0.959 0.206
Table 5: Quantitative comparison with other video editing techniques on CLIP-Image similarity and Warp Error.
Whole-body driving. In Fig. 9, we display our Qffusion can also perform whole-body animation, where UBCFasion [58] dataset is applied. For the conditions, we use DWPose [55] to detect landmarks for the face and body. Specifically, we provide the visual comparison with the current state-of-the-art methods: DreamPose [25], BDMM [57], and Animate Anyone [22]. We apply the re-produced version [23] for Animate Anyone since its official code and dataset are not publicly available. These methods are either carefully designed for whole-body animation tasks (i.e., DreamPose, BDMM), or need an additional heavy appearance network to encode the appearance of input identities (i.e., Animate Anyone). It is difficult for DreamPose to ensure clothing consistency. Besides, Animate Anyone struggles to guarantee facial fidelity. Without additional modules, our method can obtain better results than these task-specific methods, demonstrating its generalizability.
Figure 10: Jump cut smooth [51] results using our Qffusion. It has potential value for speech video editing and the film industry.
[…]
DreamPose [25] 15.21 0.803 4.688 0.165
BDMM [57] 22.26 0.853 1.516 0.094
Animate Anyone [22] 17.23 0.762 7.599 0.206
Ours 23.42 0.856 0.901 0.092
Table 6: Quantitative comparisons with the current sate-of-the-art whole-body driving methods, where our method yields the best performance on PSNR, SSIM, LPIPS, and Warp Error.
In addition to the qualitative comparison, we also give the quantitative comparison in Tab. 6. It can be seen that our method can achieve better results than those task-specific state-of-the-art techniques.
Jump Cut Smooth. A jump cut brings an abrupt, sometimes unwanted change in the viewing experience. Our method can be used to smooth these jump cuts. Fig. 10 presents an extra application that our Qffusion can deal with, i.e., jump cut smooth [51]. The application is performed as follows. 1) We take the jump-cut start frame as $\mathbf{I}^{s}$, end frame as $\mathbf{I}^{e}$. 2) We extract these two frames’ conditions $\mathbf{C}^{s}$ and $\mathbf{C}^{e}$. 3) Assuming there are $K$ frames to be generated, we interpolate intermediate conditions $\mathbf{C}^{k}_{P}=(k/K)\mathbf{C}^{s}_{P}+(1-k/K)\mathbf{C}^{e}_{P},k\in\{1,\dots,K\}$. Here $\mathbf{C}^{*}_{P}$ denotes the conditional 3D points and $\mathbf{C}^{*}$ is the visualization of $\mathbf{C}^{*}_{P}$. As seen in Fig. 10, our Qffusion can achieve seamless transitions between cuts, even in challenging cases where the talking head undergoes large-scale movement or rotation in the jump cut.
Figure 11: User study on the selected ratio. Our Qffusion outperforms the current state-of-the-art video editing methods.
5.6 User Study
We also provide a user study to compare our method with recently proposed video editing methods TokenFlow [16], Codef [39], and AnyV2V [27]. Specifically, we pose questions Q1: General Editing (GE), Q2: Temporal Consistency (TC), and Q3: Video Quality (VQ) to 30 anonymous participants on a crowd-sourcing platform, for randomly selected 12 video editing samples. We report the select ratio of four video editing methods in Fig. 11. Our Qffusion earns the highest user preference in all three aspects.
6 Conclusion
This paper presented a dual-frame-guided portrait video editing framework dubbed Qffusion, where we first modify the start and end frames with professional software (e.g., Photoshop or Meitu) and then propagate these modifications. Specifically, obeying an “animation for editing” principle, our Qffusion is trained as a general video animation model, which can be used for portrait video editing by treating the edited start and end frames as references. Specifically, we design a Quadrant-grid Arrangement (QGA) scheme for latent re-arrangement in SD, which captures spatial correspondence and temporal clues in a quadrant-grid design. Besides, stable arbitrary-length videos can be generated stably via our proposed recursive Quadrant-grid Propagation (QGP) inference. Our Qffusion serves as a foundational method, with the potential for future extension into various applications.
Supplementary Material
To ensure the reproducibility and completeness of this paper, we make the Appendix with 2 sections. Appendix A provides more details for our QGP inference pipeline. Appendix B presents more of our Qffusion’s video animation capacity. Appendix C discusses the limitations of our method.
Appendix A More Details for QGP Inference
To seek an even motion modeling, we propose a recursive influence strategy Quadrant-grid Propagation (QGP), in which we use generated frames as reference frames for next inference iteration. Specifically, given a start frame $\mathbf{I}^{s}$, an end frame $\mathbf{I}^{e}$, and a sequence of conditions, this method takes the conditions as driving signals to recursively generate intermediate frames. The pseudo-code of the proposed inference pipeline is presented at Algorithm 1, which mainly involves 4 steps:
  • Given the two reference images $\mathbf{I}^{i}$ and $\mathbf{I}^{i+k}$ ($i$, $k$ indicate the start and the interval between start and ending frame, i.e., for the first step, $i=0,k=K$), we first generate the two intermediate images $\tilde{\mathbf{I}}^{\frac{k}{3}},\tilde{\mathbf{I}}^{\frac{2k}{3}}$, where the index in Quadrant-grid Arrangement is $\left[0,\left\lfloor\frac{K}{3}\right\rfloor,\left\lfloor\frac{2K}{3}\right\rfloor,K\right]$. This makes the intervals between each sub-image in the quadrant the same, leading to a consistent temporal sampling.
  • Then we repeat step (i) to generate frames between $\left[0,\left\lfloor\frac{K}{3}\right\rfloor\right]$, $\left[\left\lfloor\frac{K}{3}\right\rfloor,\left\lfloor\frac{2K}{3}\right\rfloor\right]$, and $\left[\left\lfloor\frac{2K}{3}\right\rfloor,K\right]$ following a consistent interval $\left\lfloor\frac{k}{9}\right\rfloor$.
  • We repeat step (ii) for the overall sequence. This process effectively establishes the relationship between the generated video frames and their preceding and succeeding frames.
  • We organize the intermediate video frames to form a newly generated video.
Algorithm 1 Quadrant-grid Propagation of Qffusion
Input:
  Reference images $\mathbf{I}^{0},\mathbf{I}^{K}$, Conditions $\{\mathbf{C}^{0},\dots,\mathbf{C}^{K}\}$
Output:
  Generated new video $\{\tilde{\mathbf{I}}^{0},\dots,\tilde{\mathbf{I}}^{K}\}$
1:Queue = [[0, K]]   #[start index, interval between start and end]
2:while Queue is not empty do
3:       $i,k$ = Queue[0]
4:       Queue.popleft()
5:       $\mathcal{Q}_{0}$ = QGA ($\mathbf{I}^{i},\mathbf{C}^{i},\mathbf{C}^{i+\frac{k}{3}},\mathbf{C}^{i+\frac{2k}{3}},\mathbf{C}^{i+k},\mathbf{I}^{i+k}$)
6:       $\tilde{\mathcal{Q}}_{0}$ = Diffusion&Denoising($\mathcal{Q}_{0}$)
7:       $\tilde{\mathbf{z}}^{i+\frac{k}{3}},\tilde{\mathbf{z}}^{i+\frac{2k}{3}}=\texttt{Split\&Unstack}(\tilde{\mathcal{Q}}_{0})$
8:       $\tilde{\mathbf{I}}^{i+\frac{k}{3}},\tilde{\mathbf{I}}^{i+\frac{2k}{3}}=\texttt{Split}(\mathcal{D}([\tilde{\mathbf{z}}^{i+\frac{k}{3}};\tilde{\mathbf{z}}^{i+\frac{2k}{3}}]))$
9:       for j in [0, 1, 2] do
10:             if $i+k>K$ then
11:                    break
12:             end if
13:             Queue.append($[i+\frac{k}{3}*j,\frac{k}{3}]$)   #append processed index
14:       end for
15:end while
16:return $\{\tilde{\mathbf{I}}^{0},\dots,\tilde{\mathbf{I}}^{K}\}$
We give an example of our QGP inference strategy. Assuming there is an 82-frame video indexed as $\{0-th,1-st,...,80-th,81-st\}$ to be generated, in the first iteration, we apply the $0-th$ and $81-st$ frames as references to generate two intermediate frames: $27-th$ and $54-th$. Then, in the second iteration, we use the newly generated $27-th$ and $54-th$ as references. Concretely, we make $\{0-th,27-th\}$, $\{27-th,54-th\}$ and $\{54-th,81-st\}$ as references separately to synthesize new intermediate frames, forming $\{0-th,9-th,18-th,27-th\}$, $\{27-th,36-th,45-th,54-th\}$, and $\{54-th,63-th,72-nd,81-st\}$. The process would be stopped until all frames are produced. Note that our QGP inference gradually uses generated frames at the current inference iteration (e.g., $27-th$ and $54-th$) as reference frames for the next iteration, making the intervals between each sub-image in the four-grid representation the same in each iteration.
In contrast, the naive inference method fixs the start and end frames ($0-th$ and $81-st$) as references to gradually generate all intermediate frames $\{1-st,2-nd,...79-th,80-th\}$.
Figure 12: More qualitative comparisons with other video animation methods.
Appendix B More Results on Portrait Video Animation
Recall that in Fig. 4 of our main paper, we provide a qualitative comparison with different video animation methods, including: (i) GAN [32], (ii) ControlNet [60], and (iii) ControlNet $+$ Animatediff [18] . Here, we provide more visual comparison in Ap-Fig. 12. The results demonstrate that compared with these animation methods, our Qffusion performs the best animation in terms of ID consistency and condition alignment.
Figure 13: Our limitation. Our Qffusion struggles with ID fidelity when using different identities as reference images.
Appendix C Limitations
While our method provides promising results for various applications, there still exist some limitations. For instance, as seen in Fig. 13, our Qffusion sometimes faces unsatisfying cross-ID portrait video animation results. The reason here is that when the driving landmarks come from a different person, the shape information cannot be well-aligned between different persons.